OpenStudy (anonymous):

1+b>3 Whats the solution too the inequality.

OpenStudy (tkhunny):

Subtract one from each side.

OpenStudy (anonymous):

which side.

OpenStudy (tkhunny):

"each"? Always each. Never only one.

OpenStudy (anonymous):

how do i subtract from b.

OpenStudy (anonymous):

There is two sides to your problem. Think of > as an equals sign. 1 + b > 3 -1 - 1 b > 2

OpenStudy (tkhunny):

You don't (1 - b) - 1 = 1 - b - 1 = 1 - 1 - b = (1 - 1) - b = 0 - b = -b There is a reason why I picked "1". You do (1+b) - 1